is there a way to combine a human's speed with 6.02x10^23 to make them as fast as light?
1
Expert's answer
2019-11-11T07:48:41-0500

The relativistic mass of the moving at some velocity body can be expressed as:


"M=\frac{m_0}{\sqrt{1-\frac{v^2}{c^2}}},"


where v - the velocity of the body and m0 - the mass of the non-moving body;


We can actually express m0 via NA number:


"M=\frac{MolarMass*Number of Particles}{N_A \sqrt{1-\frac{v^2}{c^2}}},"


Now, it is clear that if the body moves at the speed close to the speed of light, then its mass M is going to be closer and closer to the infinity. And this requires an infinite amount of energy. Thus, actually, it is not possible for the body with a non-zero rest mass to move with speed of light.